# CordViP: Correspondence-based Visuomotor Policy for Dexterous Manipulation in Real-World

### v1

@@ V Conclusions and Limitations
In this paper, we present CordViP, a novel framework that learns correspondence-based visuomotor policy for dexterous manipulation in the real world. First, we leverage robust 6D pose estimation of objects and robot proprioception to obtain interaction-aware point clouds, enhancing the quality of the point clouds and addressing the issues caused by occlusions from the dexterous hand. Second, we use object-centric contact map and coordination information to design a pre-training task that effectively establishes spatial and temporal correspondences. Finally, the features obtained from the pre-trained encoder are used as conditions to train a visuomotor policy. CordViP significantly outperforms state-of-the-art 2D and 3D baselines on four real-world dexterous manipulation tasks, demonstrating highly competitive performance in both effectiveness and efficiency.
Limitations. Despite the exceptional performance demonstrated by CordViP, there are still certain limitations that could be explored in future work. First, our method struggles to accurately estimate the 6D pose of deformable objects due to the limited expressive capacity of FoundationPose. Second, the accuracy of digital twin modeling can significantly impact pose tracking and the quality of initial object point clouds, which are critical for fine-grained dexterous manipulation. One potential solution could be the inclusion of an additional wrist-mounted camera. We leave further exploration of these possibilities for future work.

### v2

@@ V Conclusions and Limitations
In this paper, we present CordViP, a novel framework that learns correspondence-based visuomotor policy for dexterous manipulation in the real world. First, we utilize powerful 3D generation methods to obtain 3D models of objects, then use robust 6D object pose estimation based on this modeling information and robot proprioception to obtain interaction-aware point clouds, enhancing the quality of the point clouds and addressing the issues caused by occlusions from the dexterous hand. Second, we use object-centric contact map and coordination information to design a pre-training task that effectively establishes spatial and temporal correspondences. Finally, the features obtained from the pre-trained encoder are used as conditions to train a visuomotor policy. CordViP significantly outperforms state-of-the-art 2D and 3D baselines on six real-world dexterous manipulation tasks, demonstrating highly competitive performance in both effectiveness and efficiency.
Limitations. Despite the exceptional performance demonstrated by CordViP, there are still certain limitations that could be explored in future work. First, our method struggles to accurately estimate the 6D pose of deformable objects due to the limited expressive capacity of FoundationPose when handling non-rigid geometry. Second, the accuracy of digital twin modeling can significantly impact pose tracking and the quality of initial object point clouds, which are critical for fine-grained dexterous manipulation. Additionally, if the dexterous hand completely occludes the object, FoundationPose may fail to track it. One promising research direction could be the incorporation of visuotactile 6D pose tracking. We leave further exploration of these possibilities for future work.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (62476011) and the Joint R&D Fund of Beijing Smart-chip Microelectronics Technology Co., Ltd (SGSC0000RGJS2401830, SGSC0000RGJS2401829).